# On the Impacts of Contexts on Repository-Level Code Generation

### v3

@@ 8 Limitations
Appendix B Coverage Enhancement Effectiveness
@@
Appendix C Test case generation
@@
Coverage Enhancement Prompts Prompt 1: Here are some Python unit test functions and the focal function that they test: # Test functions: {existing_test_functions} # Focal function: {function_under_test} Write more unit test functions that will increase the test coverage of the function under test. ——————————————————————————————————————– Prompt 2: Here are some Python unit test functions and the focal function that they test: # Test functions: {existing_test_functions} # Focal function: {function_under_test} Write more unit test functions that will cover corner cases missed by the original and will increase the test coverage of the function under test. ——————————————————————————————————————– Prompt 3: Here is a focal function under test: {function_under_test} This function under test can be tested with these Python unit test functions: {existing_test_functions} Here is an extended version of the unit test function that includes additional unit test cases that will cover methods, edge cases, corner cases, and other features of the function under test that were missed by the original unit test functions:
Appendix D Data Formating
D.1 BasePrompt
@@
D.2 InstructPrompt
Instruction Prompt Templates Prompt 1: ### Instruction: Write a Python function ‘{target_function_signature}‘ to solve the following problem: {target_function_docstring} ### Response: {BasePrompt} ——————————————————————————————————————– Prompt 2: ### Instruction: {dependency_context} The provided code snippet includes necessary dependencies for implementing the ‘{target_function_name}‘ function. Write a Python function ‘{target_function_signature}‘ to solve the following problem: {target_function_docstring} ### Response: {target_function_prompt}
@@ D.2 InstructPrompt
Appendix E Studied LLMs: Supplemental results
@@
Additionally, Table LABEL:tab:examples_gen presents examples that support our findings in Section 6.1. For instance, in the first example, we observe that the instruction-tuned model can effectively utilize the given dependencies to manage edge cases, whereas the pretrained model fails to do so. This supports our initial findings. Meanwhile, the second and third examples demonstrate that pretrained models often reimplement or devise workarounds instead of leveraging the available context. Besides, in the second example, the instruction-tuned model correctly identifies the relevant case but fails to generate the correct solution. This may suggest that hallucinations complicate the outputs generated by instruction-tuned models.
@@
Appendix F Multi-round Debugging
@@
Appendix G Context length analysis
Long-context models can enhance the ability to comprehend and select relevant context from lengthy inputs to effectively solve the required task. In this section, we examine how the context length supported by each model affects their performance in RepoExec. Figure 9 demonstrates the relationship between support context length, model size, and model family in relation to pass@k and DIR scores. The size of the dots indicates the model size, while the color represents the model family.
Figure 9: Correlation of context length to the model performance on RepoExec.
We observe that context length has a weak correlation with model performance on our dataset. In contrast, model size (scaling law) and model family, which encompass different training methods (pretraining or instruction tuning), training datasets, and architectures, show a more significant impact. The weak correlation with context length can be explained by our approach’s ability to already capture relevant information (e.g. dependency) for each data sample, resulting in the pruning of context length for practical usage (363 tokens on average shown in Table 2). Meanwhile, models that support long contexts are often trained on data containing a mix of relevant and irrelevant information and are evaluated on their ability to retrieve the correct context in a needle-in-a-haystack scenario (Roziere et al., 2023; Ivgi et al., 2023; Liu et al., 2024). Therefore, models with varying context lengths might show a weak correlation to performance in our scenario.
Appendix H Dependency extraction tool usage
We present pydepcall, a Python library designed to extract function dependencies from any repository. We provide a brief overview of its usage in the following code snippet.
@@ Appendix H Dependency extraction tool usage
# Prompt: import base64 import random import unicodedata ... class InvalidInputError(TypeError): ... def is_string(obj: Any) -> bool: ... def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ ------------------------------------------------------- # Solution: def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ if not is_string(input_string): raise InvalidInputError(input_string) return input_string[::-1] # Model type: Pretrained # Passed: No def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ return input_string[::-1] ------------------------------------------------------- # Model type: Instruction-tuned # Passed: Yes def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ if not is_string(input_string): raise InvalidInputError(input_string) return input_string[::-1]
@@
# Prompt: import random import luigi import luigi.configuration ... class hadoopcli(luigi.Config): command = ... version = ... def get_configured_hadoop_version(): """ CDH4 (hadoop 2+) has a slightly different syntax for interacting with hdfs via the command line. The default version is CDH4, but one can override this setting with "cdh3" or "apache1" in the hadoop section of the config in order to use the old syntax. """ ------------------------------------------------------- # Solution: def get_configured_hadoop_version(): """ CDH4 (hadoop 2+) has a slightly different syntax for interacting with hdfs via the command line. The default version is CDH4, but one can override this setting with "cdh3" or "apache1" in the hadoop section of the config in order to use the old syntax. """ return hadoopcli().version.lower() # Model type: Pretrained # Passed: Yes def get_configured_hdfs_client(): """ This is a helper that fetches the configuration value for ’client’ in the [hdfs] section. It will return the client that retains backwards compatibility when ’client’ isn’t configured. """ return luigi.configuration.get_config().get(’hdfs’, ’client’, ’hadoopcli’) ------------------------------------------------------- # Model type: Instruction-tuned # Passed: No def get_configured_hadoop_version(): """ CDH4 (hadoop 2+) has a slightly different syntax for interacting with hdfs via the command line. The default version is CDH4, but one can override this setting with "cdh3" or "apache1" in the hadoop section of the config in order to use the old syntax. """ hadoop_cli = hadoopcli().command.split() hadoop_version = hadoopcli().version if hadoop_version == "cdh3": return hadoop_cli + ["fs"] elif hadoop_version == "apache1": return hadoop_cli + ["dfs"] else: return hadoop_cli
@@
# Prompt: import asyncio import builtins import collections ... def is_future(x: Any) -> bool: return isinstance(x, FUTURES) def _create_future() -> Future: future = Future() # type: Future ... def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ ------------------------------------------------------- # Solution: def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ if is_future(x): return x else: fut = _create_future() fut.set_result(x) return fut # Model type: Pretrained # Passed: Yes def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ if is_future(x): return x future = Future() # type: Future future.set_result(x) return future ------------------------------------------------------- # Model type: Instruction-tuned # Passed: Yes def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ if isinstance(x, Future): return x elif isawaitable(x): return asyncio.ensure_future(x) else: future = Future() future.set_result(x) return future

### v4

@@ Appendix A Data Source
Appendix C Coverage Enhancement Effectiveness
@@
Appendix D Test case generation
Generated test quality discussion:
To ensure generated tests align with the target function’s requirements, we employ several quality control measures, as discussed in Section 4.2. First, we provide the full target function in the LLM prompt to guide the generation of tests with the intended input and output formats. We then apply a Syntax-based filter to confirm that the generated tests correctly call the function with its signature and parameters. Additionally, to verify the accuracy of the tests, we execute them to confirm their pass status and measure their line coverage.
In some minor cases, weak tests or even wrong ones may still arise despite these measures. For instance, flaky tests could pass and achieve high coverage but produce inconsistent outputs due to randomness. To address this, we execute tests multiple times and exclude those with inconsistent results. Similarly, concerns about unintended function calls are mitigated by deduplicating test cases and ensuring diversity in how functions are invoked. This process yields an average of 99 test cases per problem (Table 1), ensuring robust evaluation even if some weaker tests are included. While we do not require such extensive coverage for correctness, this approach enhances the likelihood of strong test cases identifying implementation issues effectively. Furthermore, the test enhancement module is provided with a few examples of generated tests from the previous step to enhance coverage of edge cases and corner cases to ensure specific requirements.
@@ Appendix D Test case generation
Test geneneration prompts:
Below, we present two prompts designed for generating initial unit tests and improving the quality of existing tests.
@@
Coverage Enhancement Prompts Prompt 1: Here are some Python unit test functions and the focal function that they test: # Test functions: {existing_test_functions} # Focal function: {function_under_test} Write more unit test functions that will increase the test coverage of the function under test. —————————————————————————————————– Prompt 2: Here are some Python unit test functions and the focal function that they test: # Test functions: {existing_test_functions} # Focal function: {function_under_test} Write more unit test functions that will cover corner cases missed by the original and will increase the test coverage of the function under test. —————————————————————————————————– Prompt 3: Here is a focal function under test: {function_under_test} This function under test can be tested with these Python unit test functions: {existing_test_functions} Here is an extended version of the unit test function that includes additional unit test cases that will cover methods, edge cases, corner cases, and other features of the function under test that were missed by the original unit test functions:
Appendix E Data Formating
E.1 BasePrompt
@@
E.2 InstructPrompt
Instruction Prompt Templates Prompt 1: ### Instruction: Write a Python function ‘{target_function_signature}‘ to solve the following problem: {target_function_docstring} ### Response: {BasePrompt} —————————————————————————————————– Prompt 2: ### Instruction: {dependency_context} The provided code snippet includes necessary dependencies for implementing the ‘{target_function_name}‘ function. Write a Python function ‘{target_function_signature}‘ to solve the following problem: {target_function_docstring} ### Response: {target_function_prompt}
@@ E.2 InstructPrompt
Appendix F Studied LLMs: Supplemental results
@@
Additionally, Table LABEL:tab:examples_gen presents examples that support our findings in Section 6.1. For instance, in the first example, we observe that the instruction-tuned model can effectively utilize the given dependencies to manage edge cases to raise an error, whereas the pretrained model fails to do so. This supports our first findings. Meanwhile, the second and third examples demonstrate that pretrained models try to reimplement or devise workarounds instead of leveraging the available context. Additionally, in the second example, the instruction-tuned model successfully identifies the relevant dependencies; however, it complicates the situation and fails to produce the correct solution. This may suggest that hallucinations complicate the outputs generated by instruction-tuned models. In the third example, both types of models successfully pass the tests. However, although the context provides the dependency for creating a new Future object, these models attempt to reimplement the _create_future function but fail to optimize memory usage effectively. This observation implies the potential of code smell and technical debt in these generated codes.
Besides, we present supplementary results for both closed-source and open-source models in Table 8. Specifically, we employ greedy decoding and report pass@1 and DIR metrics, selecting the best-performing results for each model based on either BasePrompt or InstructPrompt with Full context. The results demonstrate strong performance in both pass@k and DIR metrics with advanced reasoning models like DeepSeek-R1 and GPT-4o, highlighting their reliability in code generation. However, all pass@1 results remain below 50%, underscoring the challenges posed by RepoExec.
ID Model Size (B) Pass@1 DIR 1 DeepSeek-R1 Guo et al. (2025) 671 42.57 70.86 2 DeepSeek-V3 Liu et al. (2024a) 671 42.00 80.35 3 Llama 3.1-Instruct Dubey et al. (2024) 405 39.43 73.26 4 GPT-4o - 38.00 68.15 5 QwQ-Preview Team (2024) 32 37.43 57.07 6 DeepSeek-Coder-Instruct Guo et al. (2024) 33 35.71 67.02 7 Yi-1.5 Young et al. (2024) 34 35.43 61.34 8 Qwen-2.5-Coder-Instruct Hui et al. (2024) 14 34.28 65.68 9 CodeQwen1.5 Bai et al. (2023) 7 32.29 59.48 10 OpenCodeInterpreter Zheng et al. (2024) 33 31.42 61.28 11 Phi-4 Abdin et al. (2024) 14 30.86 71.10 12 GPT-4o-mini - 30.29 74.75 13 Llama 3.3-Instruct Dubey et al. (2024) 70 29.43 67.20 14 DeepSeek-Coder Guo et al. (2024) 6.7 28.86 57.44 15 Qwen-2.5-Instruct Yang et al. (2024) 14 28.85 65.63 16 starcoder2 Lozhkov et al. (2024) 15 28.57 60.97 17 Gemma2 Team et al. (2024) 27 28.00 71.68 18 Llama-3.1-Instruct Dubey et al. (2024) 70 25.14 66.95
Table 8: Performance of various LLMs on RepoExec.
@@
Appendix G Multi-round Debugging
@@
Appendix H Context length, model size and families analysis
Long-context models can enhance the ability to comprehend and select relevant context from lengthy inputs to effectively solve the required task. In this section, we examine how the context length supported by each model affects their performance in RepoExec. Figure 9 demonstrates the relationship between support context length, model size, and model family in relation to pass@k and DIR scores.
Figure 9: Correlation of context length to the model performance on RepoExec. The size of the dots indicates the model size, while the color represents the model family.
We observe that context length has a weak correlation with model performance on our dataset. In contrast, model size (scaling law) and model family, which encompass different training methods (pretraining or instruction tuning), training datasets, and architectures, show a more significant impact. The weak correlation with context length can be explained by our approach’s ability to already capture relevant information (e.g. dependency) for each data sample, resulting in the pruning of context length for practical usage (363 tokens on average shown in Table 1). Meanwhile, models that support long contexts are often trained on data containing a mix of relevant and irrelevant information and are evaluated on their ability to retrieve the correct context in a needle-in-a-haystack scenario (Roziere et al., 2023; Ivgi et al., 2023; Liu et al., 2024b). Therefore, models with varying context lengths might show a weak correlation to performance in our scenario.
Appendix I Dependency Extraction Tool
We present pydepcall, a Python library designed to extract function dependencies from any repository. Our tool provides a quick, practical solution for scalability, requiring only three lines of code to extract dependencies for all functions in any repository. Specifically, our tool is based on two conventions for retrieving function dependencies of a programming language:
  • Import Convention (for cross-file dependencies): The convention mentions how a programming language imports modules within a repository. This can be utilized to extract dependency names from different files. For instance, in Python, this can be done using statements such as "from file import function/class/variable" or "from file1.file2 import function/class/variable."
  • Calling Dependencies Convention: This principle pertains to how a program invokes variables, functions, and classes. In Python, functions are typically called by specifying their names followed by parentheses (e.g., function1(param1, param2)), or by using a dot notation to access attributes or methods from a module (e.g., file1.function1, class1.attribute1).
Besides, pydepcall can extract dependencies at various depths, up to 100 levels deep, which can help the community explore the integration of a deeper context. We provide a brief overview of its usage in the following code snippet.
@@ Appendix I Dependency Extraction Tool
# Prompt: import base64 import random import unicodedata ... class InvalidInputError(TypeError): ... def is_string(obj: Any) -> bool: ... def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ -------------------------------------------------- # Solution: def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ if not is_string(input_string): raise InvalidInputError(input_string) return input_string[::-1] # Model type: Pretrained # Passed: No def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ return input_string[::-1] -------------------------------------------------- # Model type: Instruction-tuned # Passed: Yes def reverse(input_string: str) -> str: """ Returns the string with its chars reversed. *Example:* $>$ reverse(’hello’) # returns ’olleh’ :param input_string: String to revert. :type input_string: str :return: Reversed string. """ if not is_string(input_string): raise InvalidInputError(input_string) return input_string[::-1]
@@
# Prompt: import random import luigi import luigi.configuration ... class hadoopcli(luigi.Config): command = ... version = ... def get_configured_hadoop_version(): """ CDH4 (hadoop 2+) has a slightly different syntax for interacting with hdfs via the command line. The default version is CDH4, but one can override this setting with "cdh3" or "apache1" in the hadoop section of the config in order to use the old syntax. """ -------------------------------------------------- # Solution: def get_configured_hadoop_version(): """ CDH4 (hadoop 2+) has a slightly different syntax for interacting with hdfs via the command line. The default version is CDH4, but one can override this setting with "cdh3" or "apache1" in the hadoop section of the config in order to use the old syntax. """ return hadoopcli().version.lower() # Model type: Pretrained # Passed: Yes def get_configured_hdfs_client(): """ This is a helper that fetches the configuration value for ’client’ in the [hdfs] section. It will return the client that retains backwards compatibility when ’client’ isn’t configured. """ return luigi.configuration.get_config().get(’hdfs’, ’client’, ’hadoopcli’) -------------------------------------------------- # Model type: Instruction-tuned # Passed: No def get_configured_hadoop_version(): """ CDH4 (hadoop 2+) has a slightly different syntax for interacting with hdfs via the command line. The default version is CDH4, but one can override this setting with "cdh3" or "apache1" in the hadoop section of the config in order to use the old syntax. """ hadoop_cli = hadoopcli().command.split() hadoop_version = hadoopcli().version if hadoop_version == "cdh3": return hadoop_cli + ["fs"] elif hadoop_version == "apache1": return hadoop_cli + ["dfs"] else: return hadoop_cli
@@
# Prompt: import asyncio import builtins import collections ... def is_future(x: Any) -> bool: return isinstance(x, FUTURES) def _create_future() -> Future: future = Future() # type: Future # Fixup asyncio debug info by removing extraneous stack entries source_traceback = getattr(future, "_source_traceback", ()) while source_traceback: # Each traceback entry is equivalent to a # (filename, self.lineno, self.name, self.line) tuple filename = source_traceback[-1][0] if filename == __file__: del source_traceback[-1] else: break return future def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ -------------------------------------------------- # Solution: def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ if is_future(x): return x else: fut = _create_future() fut.set_result(x) return fut # Model type: Pretrained # Passed: Yes def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ if is_future(x): return x future = Future() # type: Future future.set_result(x) return future -------------------------------------------------- # Model type: Instruction-tuned # Passed: Yes def maybe_future(x: Any) -> Future: """Converts ‘‘x‘‘ into a ‘.Future‘. If ‘‘x‘‘ is already a ‘.Future‘, it is simply returned; otherwise it is wrapped in a new ‘.Future‘. This is suitable for use as ‘‘result = yield gen.maybe_future(f())‘‘ when you don’t know whether ‘‘f()‘‘ returns a ‘.Future‘ or not. .. deprecated:: 4.3 This function only handles ‘‘Futures‘‘, not other yieldable objects. Instead of ‘maybe_future‘, check for the non-future result types you expect (often just ‘‘None‘‘), and ‘‘yield‘‘ anything unknown. """ if isinstance(x, Future): return x elif isawaitable(x): return asyncio.ensure_future(x) else: future = Future() future.set_result(x) return future